Two groups of children were given visual acity tests,group.group 1 was composed of 11 children who receive their health care from private physician.the mean score for this group was 26 with standard deviation of 5.group 2 was composed of 14 children who receive their health care from the health department and had an average score of 21 with a standard deviation of assume normally distributed population with equal variance calculate the 95%confidence interval

1
Expert's answer
2022-01-31T16:27:04-0500

A F-test is used to test for the equality of variances. The following F-ratio is obtained:


"F=\dfrac{s_1^2}{s_2^2}=\dfrac{5^2}{6^2}=0.694444"

The critical values for "df_1=11-1=10,df_2=14-1=13, \alpha=0.05" are "F_L = 0.2791," "F_U=3.2497," and since "F = 0.694444," then the null hypothesis of equal variances is not rejected.


Based on the information provided, the significance level is "\alpha = 0.05," and the degrees of freedom are "df=n_1-1+n_2-1 = 11-1+14-1=23," assuming that the population variances are equal.

Hence, it is found that the critical value for this two-tailed test for "\alpha=0.05,df=23" degrees of freedom is "t_c=2.068658."

Since the population variances are assumed to be equal, we need to compute the pooled standard deviation, as follows:


"s_p=\sqrt{\dfrac{(n_1-1)s_1^2+(n_2-1)s_2^2}{n_1+n_2-2}}"

"=\sqrt{\dfrac{(11-1)5^2+(14-1)6^2}{11+14-2}}"

"=5.587253"

Since we assume that the population variances are equal, the standard error is computed as follows:


"se=s_p\sqrt{\dfrac{1}{n_1}+\dfrac{1}{n_2}}=5.587253\sqrt{\dfrac{1}{11}+\dfrac{1}{14}}"

"=2.251168"

Now, we finally compute the confidence interval:


"CI=(\bar{x_1}-\bar{x_2}-t_c\times se, \bar{x_1}-\bar{x_2}+t_c\times se)"

"=(26-21-2.068658\times 2.251168,"

"26-21+2.068658\times 2.251168)"

"=(0.3431,9.6569)"

Therefore, based on the data provided, the 95% confidence interval for the difference between the population means "\mu_1 - \mu_2" is "0.343 1< \mu_1 - \mu_2 < 9.6569," which indicates that we are 95% confident that the true difference between population means is contained by the interval "(0.3431, 9.6569)."